find the dimensions of the rectangle. if its length is twice its width. perimeter is equal to 208.
Answer by checkley77(12844)

L=2W
2*W+(2*2W)=208
2W+4W=208
6W=208
W=208/6
W=34.66667 ANSWER FOR THE WIDTH.
L=2*34.66667
L=69.33333 ANSWER.
PROOF:
2*34.6667+2*2*34.6667=208
69.3333+138.66667=208
208=208